# Comprehend, Divide, and Conquer: Feature Subspace Exploration via Multi-Agent Hierarchical Reinforcement Learning

### v1

@@ 5.2. Space/Time Complexity Analysis
Space/Time Complexity Comparison with Baselines. This experiment aims to answer the question: Does HRLFS excel in both time and performance? Figure 6 shows the performance and time consumption of HRLFS compared to other baseline methods and our ablation model $\textsc{HRLFS}^{-h}$ which replaces the multi-agent hierarchical reinforcement learning architecture with One-Agent-Per-Feature (OAPF) architecture. We found that HRLFS outperforms all baselines and $\textsc{HRLFS}^{-h}$, with time consumption comparable to traditional statistical methods (KBest, MCDM, and LASSONet) and much superior to methods based on RL (SARLFS and MARLFS), deep learning-based methods (SAFS, CoFS, and GAINS), and $\textsc{HRLFS}^{-h}$. Compared to the deep learning-based methods, our agents only consists of two-layer neural networks, which are less time-consuming compared to numerous-parameters models used by these methods. Compared to RL-based methods, HRLFS applies the hierarchical decision architecture to manage clustered features, greatly reducing the number of active agents, where active agents will make decisions when the rest of the agents remain inactive, enabling more efficient iteration and less time consumption. Further, HRLFS far outperforms $\textsc{HRLFS}^{-h}$ in both performance and time, underlining the fact that the hierarchical decision architecture can capture correlations between features, facilitate agents’ collaboration, and reduce the number of active agents. In conclusion, with effective state extraction and hierarchical architecture design, HRLFS offers great advantages in terms of time and performance.
@@ 5.5. Analysis of the Hybrid State Extraction Method
We compared various feature distribution extraction methods and semantic extraction models with different parameter scales in several datasets, shown in Figure 10. Statistic+GPT-4 applies the mean, maximum, minimum, variance, the first quartile, the second quartile, and the third quartile to extract feature distribution and GPT-4 to leverage metadata. GMM+BERT, GMM+QWen, GMM+Gemini use GMM to pick up statistic and leverage semantic information using BERT (bert, ), gte-Qwen2-7B-instruct (gte-Qwen2-7B-instruct, ) (called via Python API dashscope.TextEmbedding.call), gemini-embedding-exp  (gemini-embedding-exp-03-07, ) (called via Python API google.Client.models.embed_content) respectively. By comparing Statistic+GPT-4 and GMM+GPT-4, we can observe that feature distribution obtained using GMM has better results compared to statistics. The underlying driver of this behavior is that statistics such as mean, maximum, minimum, etc. can only provide a simple description of the data and cannot capture the inner structure of the data, whereas GMM can approximate an arbitrarily shaped probability distribution by a linear combination of multiple Gaussian distributions, which better captures this complexity and leads to better results. By comparing GMM+BERT, GMM+QWen, GMM+Gemini, GMM+GPT-4, we find that GPT-4 extracts semantic information with better performance than other models. This implies that compared to BERT and QWen, the model architecture and scale of GPT-4 is larger and capable of capturing more complex semantic relationships and contextual information. Compared to Gemini, GPT-4’s data screening and cleaning criteria reduce errors and biases in training data, further improving the accuracy and reliability of semantic information extraction. In summary, as GMM and GPT-4 can accurately extract feature distribution and semantic information from data, they are the best choice for feature state extraction.
@@ 5.6. Analysis of the Feature Clustering Algorithm Selection
5.7. Analysis of Reinforcement Learning Backend Selection
This experiment aims to answer the question: Whether the policy model used in HRLFS is superior to other existing models?
@@ 5.7. Analysis of Reinforcement Learning Backend Selection
(b) Multi Classification Task
@@
We compared ActorCritic (haarnoja2018soft, ) (Ours) with DQN (dqn, ), Double DQN (double_dqn, ) and Dueling DQN (dueling_dqn, ) on binary classification, multi classification and regression tasks in Figure 12. We find that ActorCritic outperforms other methods in all tasks. This implies that, compared to methods such as DQN, which rely on indirect updating of Q-values, ActorCritic combines policy gradient and value function estimation and is able to directly optimize the policy, reducing variance and improving learning efficiency, resulting in better performance on all tasks. In conclusion, ActorCritic is a suitable choice for the policy model due to its higher efficiency and stability of strategy optimization.
@@ 7. Conclusion and Remark
This paper studies the challenges of exploring feature subspaces from complex datasets. Specifically, we reformulated the problem of feature selection as a multi-agent hierarchical reinforcement learning framework through the hybrid feature state extraction (comprehend), the hierarchical agent initialization (divide), and the exploration and optimization (conquer). We carried out comprehensive evaluations of HRLFS, illusrating its efficient, superior, and robust performance in various data sets from different tasks and fields.

### v2

@@ 5.2. Space/Time Complexity Analysis
Space/Time Complexity Comparison with Baselines. This experiment aims to answer the question: Does HRLFS excel in both time and performance? Figure 6 shows the performance and time consumption of HRLFS compared to other baseline methods and our ablation model $\textsc{HRLFS}^{-h}$ which replaces the multi-agent hierarchical reinforcement learning architecture with One-Agent-Per-Feature (OAPF) architecture. We found that HRLFS outperforms all baselines and $\textsc{HRLFS}^{-h}$, with time consumption comparable to traditional statistical methods (KBest, MCDM, and LASSONet) and much superior to methods based on RL (RLAS and MARLFS), deep learning-based methods (SAFS, CoFS, and GAINS), and $\textsc{HRLFS}^{-h}$. Compared to the deep learning-based methods, our agents only consists of two-layer neural networks, which are less time-consuming compared to numerous-parameters models used by these methods. Compared to RL-based methods, HRLFS applies the hierarchical decision architecture to manage clustered features, greatly reducing the number of active agents, where active agents will make decisions when the rest of the agents remain inactive, enabling more efficient iteration and less time consumption. Further, HRLFS far outperforms $\textsc{HRLFS}^{-h}$ in both performance and time, underlining the fact that the hierarchical decision architecture can capture correlations between features, facilitate agents’ collaboration, and reduce the number of active agents. In conclusion, with effective state extraction and hierarchical architecture design, HRLFS offers great advantages in terms of time and performance.
@@ 5.5. Analysis of the Hybrid State Extraction Method
We compared various feature distribution extraction methods and semantic extraction models with different parameter scales in several datasets, shown in Figure 10. Statistic+GPT-4 applies the mean, maximum, minimum, variance, the first quartile, the second quartile, and the third quartile to extract feature distribution and GPT-4 to leverage metadata. GMM+BERT, GMM+QWen, GMM+Gemini use GMM to pick up statistic and leverage semantic information using BERT (bert, ), gte-Qwen2-7B-instruct (gte-Qwen2-7B-instruct, ), gemini-embedding-exp  (gemini-embedding-exp-03-07, ), respectively. By comparing Statistic+GPT-4 and GMM+GPT-4, we can observe that the feature distribution obtained using GMM has better results compared to statistics. The underlying driver of this behavior is that statistics such as mean, maximum, minimum, etc., can only provide a simple description of the data and cannot capture the inner structure of the data, whereas GMM can approximate an arbitrarily shaped probability distribution by a linear combination of multiple Gaussian distributions, which better captures this complexity and leads to better results. By comparing GMM+BERT, GMM+QWen, GMM+Gemini, GMM+GPT-4, we find that GPT-4 extracts semantic information with better performance than other models. This implies that compared to BERT and QWen, the model architecture and scale of GPT-4 are larger and capable of capturing more complex semantic relationships and contextual information. Compared to Gemini, GPT-4’s data screening and cleaning criteria reduce errors and biases in training data, further improving the accuracy and reliability of semantic information extraction. In summary, as GMM and GPT-4 can accurately extract feature distribution and semantic information from data, they are the best choice for feature state extraction.
@@ 5.6. Analysis of the Feature Clustering Algorithm Selection
(b) Multi-class Classification Task
@@
5.7. Analysis of Reinforcement Learning Backend Selection
This experiment aims to answer the question: Whether the policy model used in HRLFS is superior to other existing models? We compared ActorCritic (haarnoja2018soft, ) (Ours) with DQN (dqn, ), Double DQN (double_dqn, ) and Dueling DQN (dueling_dqn, ) on binary classification, multi classification and regression tasks in Figure 12. We find that ActorCritic outperforms other methods in all tasks. This implies that, compared to methods such as DQN, which rely on indirect updating of Q-values, ActorCritic combines policy gradient and value function estimation and is able to directly optimize the policy, reducing variance and improving learning efficiency, resulting in better performance on all tasks. In conclusion, ActorCritic is a suitable choice for the policy model due to its higher efficiency and stability of strategy optimization.
@@ 5.8. Hyperparameter Study
5.9. Empirical Validation of the Hierarchical Assumption
(a) Balance Factor.
(b) Total Tree Height.
Figure 14. Empirical study of HRLFS’s hierarchical structure compared with the perfect binary tree.
Our theoretical analysis assumes that the constructed hierarchical structure approximates a perfect binary tree. To assess this assumption empirically, we evaluated the balance factor and height of the learned hierarchies across all 21 datasets. The balance factor is defined as the average difference in height between the left and right subtrees for all nodes (larsen1994avl, ). In a perfect binary tree, this difference never exceeds 1, and the tree height is uniquely determined by the number of nodes. As shown in Figure 14, the absolute balance factor remains below 1 for all datasets, and over 80% of them fall within 0.5. In addition, for 18 datasets, the height of our constructed hierarchy differs from that of the corresponding perfect binary tree by no more than 2 levels. This strong similarity arises because our Agent Hierarchy Construction algorithm clusters agent sets with comparable sizes at earlier stages, naturally promoting balanced splits. In summary, both balance and height analyses confirm that the constructed hierarchies empirically resemble perfect binary trees, thereby supporting the validity of our assumption.
@@ 6. Related Work
RL-based Feature Selection. Reinforcement learning has been widely employed to recast wrapper-based feature selection as an objective-directed sequential decision problem, enabling iterative subspace exploration. In the early stage, Liu et al. (sarlfs, ) proposed a single-agent-all-feature framework; however, it suffers from poor scalability and high time complexity ($\mathcal{O}(N)$ on high-dimensional data). Liu et al. (marlfs, ) introduced a multi-agent reinforcement learning (MARL) architectures that assign one agent per feature, but the resulting quadratic growth in agent count quickly becomes prohibitive. Group-based and interaction-wise agents(fan2020autofs, ; fan2021autogfs, ) reduce the decision horizon by clustering features with simple statistical criteria, yet they ignore semantic and contextual relationships, leading to redundant or sub-optimal selections (li2024exploring, ; moraffah2024causal, ). Inspired by hierarchical RL’s success in long-horizon tasks (hengst2011hierarchical, ; WangWYKP24, ), we adopt a divide-and-conquer multi-agent hierarchical RL framework that first semantically clusters features and then assigns cooperative hierarchical agents to each cluster, simultaneously improving scalability, efficiency, and downstream performance.
@@ 7. Conclusion and Remark
This paper studies the challenges of exploring feature subspaces from complex datasets. Specifically, we reformulated the problem of feature selection as a multi-agent hierarchical reinforcement learning framework through the hybrid feature state extraction (comprehend), the hierarchical agent initialization (divide), and the exploration and optimization (conquer). We carried out comprehensive evaluations of HRLFS, illustrating its efficient, superior, and robust performance in various data sets from different tasks and fields.
8. Limitation and Future Work
HRLFS relies on metadata and feature names to generate feature descriptions, which makes it vulnerable to adversarial or maliciously crafted inputs that could corrupt the generated descriptions. Moreover, the current design and evaluation are limited to tabular data. Extending the framework to high-dimensional modalities such as time series, images, or audio remains an open challenge. Finally, the present formulation focuses on standard supervised learning tasks (classification and regression). Adapting HRLFS to more diverse downstream objectives—including survival analysis, anomaly detection, and causal inference—represents a promising direction for future research.